OpenStudy (anonymous):

why is sin^2(67.5˚) = (2√2+1)/(2√2) and not (2√2-1)/(2√2)? my working: sin^2(67.5˚) =(cos2A)/2 +1 =[cos(67.5)]/2 +1 =(cos135)/2 +1 =(-1/√2)/2 +1 =(2√2-1)/(2√2) could someone figure out what went wrong? ><

OpenStudy (anonymous):

cos(2x)=1-2sin^2(x) -1+cos(2x)=-2sin^2(x) (1-cos(2x))/2=sin^2(x) <---------------- you did not divide the one (1-cos(135)/2=sin^2(67.5) (1+1/sqrt(2))/2=sin^2(67.5) (1+sqrt(2)/2)/2=sin^2(67.5) (sqrt(2)+1)/(2sqrt(2))=sin^2(67.5)

OpenStudy (anonymous):

ohhhhh thank you for typing all that out so clearly :3